OpenStudy (anonymous):

Two runners leave from the same location traveling in opposite directions. The first runner is running at 7 mph and the second is running at 9 mph. How long have they been running when they are 24 miles apart?

OpenStudy (anonymous):

I got 1.5 hours.

OpenStudy (anonymous):

Is it right?

OpenStudy (anonymous):

r*t=d The sum of their distances is 24 miles. 7t + 9t = 24\[t=\frac{3}{2}=1.5 \text{ hours} \]\[7*\frac{3}{2}+9*\frac{3}{2}=24\text{ miles}\]